### v1

# Token Is All You Need: Cognitive Planning through Sparse Intent Alignment

Abstract
We challenge the long-standing assumption that exhaustive scene modeling is required for high-performance end-to-end autonomous driving (E2EAD). Unlike world-model approaches that rely on computationally intensive future scene generation or vision-language-action (VLA) systems constrained by Markov assumptions, we show that a minimal set of semantically rich tokens is sufficient for effective planning. Experiments on nuPlan (720 scenarios, 11k+ samples) with perception-informed BEV representations yield three key findings: (1) even without future prediction, our sparse representation achieves 0.548 m ADE—comparable to or surpassing prior methods reporting $\sim$0.75 m on nuScenes; (2) conditioning trajectory decoding on predicted future tokens reduces ADE to 0.479 m, a 12.6% improvement over current-state baselines; and (3) explicit reconstruction loss offers no benefit and may degrade performance under reliable perception inputs. Notably, we observe the emergence of temporal fuzziness: the model learns to attend adaptively to task-relevant semantics rather than aligning rigidly to fixed timestamps, providing a cognitive advantage for planning under uncertainty. Our “token is all you need” principle marks a paradigm shift—from reconstructing the world to understanding it—laying a foundation for cognitively inspired systems that plan through imagination rather than reaction.
I Introduction
End-to-end autonomous driving (E2EAD) seeks to map perception directly to control, bypassing traditional pipelines that decompose perception, prediction, and planning [1, 2, 3]. Recent works such as SSR [4] achieve strong performance through sparse BEV representations and temporal self-supervision. Yet, most approaches still assume that effective planning requires dense scene representations and future self-supervision, leading to complex architectures that simulate full scene dynamics or rely on restrictive Markov assumptions.
World-model methods attempt to generate future scenes in latent space, but incur high computational cost and error accumulation in long-horizon reasoning. Likewise, vision-language-action (VLA) frameworks [5] introduce symbolic reasoning but typically assume temporal independence, limiting their ability to handle multi-modal or uncertain futures.
Humans, by contrast, neither reconstruct every visual detail nor act solely on the present. They maintain compact cognitive models, selectively attending to semantically relevant cues while mentally simulating possible futures. Inspired by this, we hypothesize that high-performance planning arises not from reconstructing the world but from understanding it—through a minimal set of semantically meaningful tokens encoding navigational intent.
We propose the Tokenized Intent World Model (TIWM), a cognitively inspired framework that redefines future reasoning in E2EAD:
  • Sparse intent representation: environmental semantics are compressed into a small set of high-dimensional tokens capturing task-relevant information, yielding efficiency and interpretability;
  • Task-driven semantic alignment: the model aligns intent tokens directly with planning objectives, eliminating the need for pixel-level reconstruction;
  • Temporal fuzziness as a cognitive feature: reasoning probabilistically about future states enables robust and flexible planning under uncertainty.
Experiments on nuPlan validate these principles. Even without future prediction, our sparse model achieves 0.548 m ADE, challenging the belief that dense modeling is essential. Conditioning trajectory decoding on predicted future tokens further improves ADE to 0.479 m—a 12.6% gain within the same framework—demonstrating that performance emerges from semantic imagination rather than pixel reconstruction. Adding explicit reconstruction loss provides no benefit, confirming that task-driven semantic alignment alone suffices under reliable perception.
Together, these results reveal a paradigm shift: effective planning stems not from exhaustive reconstruction but from selective understanding. By treating intent as the atomic unit of world representation, TIWM bridges two previously disjoint paradigms—world models that simulate and VLA systems that interpret—showing that belief–intent co-evolution, rather than dense reconstruction, underlies human-like planning.
II Related Work
Autonomous driving has evolved from modular pipelines toward end-to-end (E2E) frameworks. Early systems decomposed perception, prediction, and planning into separate stages—interpretable but prone to error propagation and hand-crafted heuristics [6, 7]. Modern E2E approaches employ deep neural networks to map sensory inputs directly to control signals, enabling joint optimization across the full decision-making stack [1, 2, 3, 4].
Vision-based methods such as UniAD [1] and VAD [2] achieve strong results via multi-camera fusion and multi-task learning. However, their reliance on dense feature reconstruction and short-horizon prediction limits long-term reasoning and scalability. In contrast, our work demonstrates that high-performance planning can emerge from a minimal set of semantically rich tokens—without dense reconstruction or auxiliary supervision.
Latent world modeling frameworks such as World Models [8] and Dreamer [9] enable ”mental simulation” through compact latent dynamics. Yet, they face two major limitations in real-world driving: (1) the computational burden of generating full future scenes, and (2) the absence of theoretical guarantees for modeling multimodal behaviors, which leads to compounding errors in closed-loop control [10]. Our experiments show that this reconstruction is unnecessary—aligning sparse, task-relevant semantics with future intents suffices for high-quality planning. Unlike these scene-generative approaches, we focus solely on decision-critical information, showing that “token is all you need” for world modeling.
Vision-language-action (VLA) frameworks unify perception, reasoning, and control [5], often conditioning on natural-language commands to enhance interpretability. However, most assume temporal independence between observations, producing reactive rather than anticipatory policies.
Sparse scene modeling offers an alternative. SSR [4] extracts 16 navigation-guided tokens from BEV features and achieves 0.75 m ADE on nuScenes, using a Future Feature Predictor to align predicted and real BEV features via L2 loss. While conceptually aligned with our approach, SSR still relies on dense BEV feature reconstruction for self-supervision. As its authors note, “we supervise directly using an L2 loss with the real future BEV feature,” an assumption our results refute. We show that reconstruction loss not only fails to improve performance but can impede the emergence of high-level intent representations, whereas task-driven semantic alignment yields superior results.
Most E2E planners—whether perception-to-trajectory regressors or VLM/VLA hybrids [11, 12]—remain constrained by Markovian dynamics, neglecting temporal dependencies and structural similarities in high-dimensional state spaces. This often leads to multi-objective conflicts and degraded safety generalization [10]. Consequently, such systems lack genuine “mental simulation,” i.e., the capacity to anticipate multiple futures before acting. Our method introduces autoregressive intent chains that enable future-oriented reasoning through task-driven semantic alignment rather than pixel-perfect reconstruction, capturing the essence of cognitive planning while avoiding over-constrained temporal alignment.
Drawing from cognitive science, prior studies have explored attention, memory, and imagination in autonomous decision-making [13, 14, 15]. However, most remain conceptual or limited to architectural analogies. We instead operationalize two core cognitive principles—mental simulation [16, 17] and sparse coding [18]—within a unified generative–planning framework. This integration yields a computationally efficient yet cognitively grounded planner that achieves state-of-the-art performance.
In summary, while existing E2E and world-model approaches either reconstruct dense futures or assume Markovian independence, our Tokenized Intent World Model (TIWM) demonstrates that effective planning can emerge from sparse, semantically aligned intent tokens. This reframes the goal of autonomous reasoning—from reconstructing the world to understanding it—enabling agents that plan through imagination rather than reaction.
III Method
Figure 1: Tokenized Intent World Model: from perception to cognitive world. At each timestep, sparse tokens are extracted from BEV perception features, representing the distilled semantics of the scene. The model then autoregressively predicts future intent tokens—compact, task-relevant abstractions of the agent’s imagined goals. The planning decoder jointly reasons over current sparse tokens and predicted future intents to generate executable trajectories. Each intent token serves three roles: (1) Representation — compressing the core cognitive state of the environment; (2) Generation — unfolding plausible futures through autoregressive prediction; (3) Planning — directly guiding the trajectory decoder toward goal-directed action. In this framework, the world becomes an actionable cognitive entity—interpreted, imagined, and utilized by the decision system—rather than a dense intermediate reconstruction.
We propose the Tokenized Intent World Model (TIWM), a cognitively inspired planning framework that departs from conventional end-to-end autonomous driving (E2EAD) pipelines (Fig. 1). Unlike prior methods that rely on dense feature maps or pixel-level future reconstruction, TIWM achieves high-performance planning with a compact set of semantically rich tokens, requiring neither full-scene generation nor heavy self-supervision.
III-A Sparse Intent Representation
TIWM operates on perception-informed Bird’s Eye View (BEV) representations, structured as a grid with two channels for dynamic objects and four for map layers (LANE, INTERSECTION, STOP_LINE, CROSSWALK). Instead of raw image inputs requiring depth estimation, TIWM leverages reliable perception modules that provide geometric and semantic stability suitable for real-world deployment. Since our goal is to study planning rather than perception, we assume access to reliable BEV inputs to isolate decision-level behavior.
The core component is the Sparse Token Learner, which compresses a dense BEV tensor $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ into a compact set of N=16N{=}16 tokens $\mathbf{T}\in\mathbb{R}^{N\times D}$:
$\mathbf{T}=\sum_{hw}\text{Softmax}(\phi(\mathbf{X}))\odot\mathbf{X},$ (1)
where attention weights are computed via a lightweight CNN:
$\phi(\mathbf{X})=\text{Conv}_{2}(\text{ReLU}(\text{Conv}_{1}(\mathbf{X}))).$ (2)
This process condenses a $256{\times}256$ BEV grid into 16 high-level semantic tokens, each representing a critical aspect of the driving scene. This follows the cognitive principle of sparse coding, where minimal active units capture the most relevant information for decision-making.
III-B Intent Chain Evolution and Task-Driven Alignment
Instead of explicitly predicting future states, TIWM models the evolution of intent chains through an autoregressive Transformer decoder:
$\mathbf{I}_{t+1}=\text{TransformerDecoder}(\mathbf{Q},\mathbf{T}_{t}),$ (3)
where $\mathbf{T}_{t}$ are current scene tokens, $\mathbf{Q}$ are learnable queries, and $\mathbf{I}_{t+1}$ is the predicted future intent. The decoder consists of two layers with eight attention heads and an embedding dimension of D=256D{=}256.
In contrast to SSR [4], which enforces dense alignment between predicted and real BEV features, TIWM learns task-driven semantic alignment. The training objective focuses purely on the planning outcome:
$\mathcal{L}_{\text{traj}}=\lambda_{\text{traj}}\|\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1})-\mathbf{Y}_{t+1}\|_{2},$ (4)
and the overall objective is simply:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}}.$ (5)
This minimalist formulation allows the model to adaptively focus on decision-critical semantics instead of rigid temporal matching. Empirically, it yields temporal fuzziness—an emergent behavior where the model flexibly attends to semantically meaningful elements across time, improving robustness under uncertainty.
For comparison, we also evaluate an auxiliary reconstruction loss:
$\mathcal{L}_{\text{intent}}=\lambda_{\text{intent}}\|\mathbf{I}_{t+1}-\text{SparseExtractor}(\mathbf{Y}_{t+1})\|_{2},$ (6)
with a combined objective:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}}+\mathcal{L}_{\text{intent}}.$ (7)
However, this explicit alignment consistently degraded performance by over-regularizing the token space, confirming that trajectory supervision alone is sufficient to learn meaningful intent representations. This supports our core principle: “token is all you need”—semantic alignment, not reconstruction, drives effective planning.
The resulting flexibility suggests that cognitive planning benefits from probabilistic, rather than deterministic, representations of future semantics. While this work focuses on a single intent chain, extending the model to multi-intent ensembles for representing alternative futures remains an exciting future direction.
III-C Cognitive Planning Mechanism
The trajectory decoder generates motion plans conditioned on both current scene tokens and predicted future intents:
$\mathbf{Y}_{t+1:t+H}=\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1}),$ (8)
where $\mathbf{Y}_{t+1:t+H}$ denotes a 3-second trajectory over H=6H{=}6 future steps. The Multi-Modal Trajectory Predictor operates as follows:
  • Mode Selection: Predict mixture weights across $M$ driving modes based on the intent embedding:
    $\mathbf{w}=\text{softmax}(\mathbf{W}_{m}\mathbf{I}_{t+1}).$ (9)
  • Cross-Attention Fusion: Fuse current context and intent via multi-head attention:
    $\mathbf{Q}_{\text{fused}}=\text{LayerNorm}(\text{MultiHeadAttn}(\mathbf{Q},\mathbf{T}_{t},\mathbf{I}_{t+1})+\mathbf{Q}).$ (10)
  • Trajectory Generation: Decode fused queries into 2D trajectory points:
    $\mathbf{Y}=\text{MLP}(\mathbf{Q}_{\text{fused}}).$ (11)
This process embodies the cognitive notion of planning through imagination. Rather than reacting to the current state, the model anticipates possible futures through predicted intent tokens, evaluates alternatives, and plans actions accordingly. This yields a cognitively interpretable, foresight-driven decision mechanism. Empirically, removing explicit reconstruction loss ($\lambda_{\text{align}}=0$) improves both stability and accuracy, validating that task-driven semantic alignment alone enables robust planning—further reinforcing our guiding principle: “token is all you need.”
IV Experiments
IV-A Experimental Setup
Dataset. We evaluate the proposed Tokenized Intent World Model (TIWM) on the nuPlan benchmark for urban autonomous driving. Following standard practice, we use 720 scenarios ($\approx$11k samples) with a 9:1 train–validation split, covering diverse settings such as intersections, merges, and multi-agent interactions.
Input Representation. We employ NuPlan’s RasterFeatureBuilder to construct Bird’s-Eye-View (BEV) inputs from structured perception and HD map data. Four semantic layers (LANE, INTERSECTION, STOP_LINE, CROSSWALK) and two vehicle-occupancy channels are included, providing clean geometry aligned with real-world deployment where perception modules are reliable.
Evaluation Metrics. Performance is measured by Average Displacement Error (ADE) and Final Displacement Error (FDE) over a 3 s horizon at 0.5 s intervals, following the standard AVG protocol used in VAD and SSR. ADE reflects mean trajectory deviation, while FDE captures the terminal-point error.
Implementation Details. TIWM adopts a ResNet-18 encoder (d=256d{=}256) and processes four historical frames (2 s) to predict six future points (3 s). Training uses AdamW ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$), batch size 32, learning rate $1{\times}10^{-4}$ with cosine decay, and weight decay $1{\times}10^{-4}$. All models are trained for 50 epochs on a single NVIDIA RTX 4090 (24 GB).
IV-B Main Results
Table IV-B summarizes our ablation results on nuPlan, along with representative state-of-the-art (SOTA) methods on nuScenes. TIWM achieves strong performance through task-driven semantic alignment, without dense reconstruction or auxiliary supervision.
Table 1: Ablation results on nuPlan, with SOTA methods on nuScenes shown for reference. Our framework achieves high accuracy via sparse intent alignment, without explicit reconstruction. Numerical comparisons across datasets are indicative only. Method Future Intent Self-Supervision Type ADE (m) FDE (m) Params (M) Methods evaluated on nuScenes (reference only) UniAD [1] $\times$ Multi-task Auxiliary 1.03 1.65 $\sim$80 VAD-Base [2] $\times$ Multi-task Auxiliary 1.22 1.98 $\sim$55 SSR [4] $\times$ Dense Reconstruction 0.75 1.36 $\sim$32 Our ablation study on nuPlan (main results) Current token, no intent loss $\times$ None 0.548 0.993 14.2 Current token, with intent loss $\times$ Sparse Intent Reconstruction 0.554 1.007 14.2 Future token, no intent loss Task-driven Semantic Alignment 0.479 0.864 14.2 Future token, with intent loss Sparse Intent Reconstruction 0.489 0.890 14.2
Note: Parameter counts for SSR, UniAD, and VAD are estimated since exact values are not reported. SSR uses dense BEV reconstruction (“we supervise directly using an L2 loss with the real future BEV feature” [4]), whereas TIWM aligns only 16 sparse intent tokens—conceptually distinct from feature-level reconstruction. Within nuPlan, the Future token, no intent loss setting attains 0.479 m ADE—13.5% better than its reconstruction-based variant with current token.
Key Findings.
(1) Minimal complexity yields strong performance. Even without future prediction, TIWM achieves 0.548 m ADE, comparable to or surpassing prior benchmarks on nuScenes. This demonstrates that semantically reliable inputs enable high-quality planning without dense reconstruction. Within the same setup, removing future prediction increases ADE by only 0.006 m, showing that additional complexity contributes little.
(2) Future-conditioned decoding enhances reasoning. Conditioning trajectory generation on predicted future tokens—rather than reconstructed BEV features—reduces ADE to 0.479 m, a 12.6% gain over the baseline. This validates that planning via semantic imagination, rather than pixel-level reconstruction, drives performance.
(3) Explicit reconstruction loss hinders learning. In contrast to SSR’s reconstruction-based self-supervision, adding an intent-alignment term slightly degrades performance (0.489 m vs. 0.479 m ADE) and slows convergence (best epoch 43 vs. 40). Trajectory supervision alone provides a cleaner and more effective optimization signal.
Interpretation.
The key distinction between TIWM and SSR lies in their self-supervision philosophy. SSR [4] enforces pixel-level alignment by reconstructing future BEV features under auxiliary trajectory guidance, effectively coupling perception, prediction, and planning through dense supervision. In contrast, TIWM adopts a task-driven, token-level semantic alignment strategy—reconstructing only the compact intent representation that is directly optimized for the planning objective. This design decouples low-level feature reconstruction from high-level decision semantics, resulting in a more efficient and cognitively interpretable learning process.
IV-C Ablation Study
To assess each component’s role, we evaluate four controlled variants:
  • Current tokens, no intent loss: decode from current tokens only, without future prediction or reconstruction.
  • Current tokens, with intent loss: decode from current tokens with auxiliary reconstruction supervision.
  • Future tokens, no intent loss: decode from current and predicted future tokens without reconstruction.
  • Future tokens, with intent loss: decode from current and predicted future tokens with explicit L2 reconstruction.
Fig. 2 reports validation ADE and training loss across configurations. All models improve steadily up to epoch 50, with the best performance typically around epochs 40–45, suggesting that longer training enhances temporal reasoning.
Figure 2: Training curves for four configurations: validation ADE (left, lower is better) and training loss (right). “No intent loss” variants consistently outperform their reconstruction-based counterparts after epoch 20.
The contrast between Future tokens, no intent loss and Future tokens, with intent loss is particularly revealing. Both condition the decoder on predicted intents, yet only the former allows trajectory supervision to structure the latent token space. The reconstruction loss biases tokens toward low-level similarity, suppressing the emergence of high-level semantic alignment essential for effective planning.
IV-D Discussion
Our experiments reveal three core principles of cognitive planning:
(1) Task-driven semantic alignment surpasses pixel-level reconstruction. The consistent superiority of the No Intent Loss configuration demonstrates that semantic alignment driven solely by the trajectory objective outperforms explicit reconstruction. This aligns with human cognition, where attention selectively prioritizes task-relevant cues rather than exhaustively reconstructing the visual scene. As Li et al. [4] noted, reconstruction-heavy paradigms often require multi-objective optimization across perception, prediction, and planning, leading to unstable convergence. Similarly, Zheng et al. [10] observed that overly entangled objectives hinder error recovery and safety generalization in autonomous driving. By focusing learning on decision-critical semantics, TIWM achieves greater stability, efficiency, and interpretability.
(2) “Temporal fuzziness” is a feature, not a flaw. The model learns what to focus on rather than when, maintaining probabilistic attention over key semantic elements. This flexibility enables robust reasoning under uncertainty and avoids overfitting to strict temporal alignment—an inherent limitation of frame-based world models.
(3) Planning through imagination outperforms reactive policies. The 12.6% performance gain from future-conditioned decoding confirms that intelligence arises from anticipating multiple futures before acting. This addresses the Markovian limitation of VLA systems [12], demonstrating that intent evolution provides a computational form of mental simulation that supports foresightful planning.
Together, these insights mark a shift from reconstructing the world to understanding the world. Instead of expending computation on dense reconstruction—as in SSR [19]—TIWM achieves stronger results through lightweight, task-driven semantic alignment. This bridges two long-divided paradigms in embodied AI: world models emphasizing exhaustive scene synthesis and VLA systems limited to instantaneous decision-making. We show that belief–intent co-evolution, not pixel fidelity, underpins effective planning.
The implications of this work extend beyond autonomous driving. The “token is all you need” principle offers a general cognitive lens for embodied intelligence. In robotic manipulation, sparse intent tokens could represent affordances rather than full object geometry; in conversational AI, they could encode high-level discourse states instead of exhaustive linguistic context—focusing computation on the semantics that truly drive behavior.
More broadly, these findings suggest that intelligence does not arise from reconstructing every detail of the environment, but from learning to capture and reason over the information that matters most for the task at hand. TIWM provides not only an efficient planning framework but also an initial step toward unifying representation, intent, and prediction—a direction consistent with cognitive theories of internal world modeling. By grounding decision-making in tokenized, semantically aligned representations, TIWM bridges the gap between data-driven perception and cognitively inspired reasoning, pointing toward a new generation of agents that plan through understanding rather than reaction.
V Conclusion
We challenge the prevailing assumption that high-fidelity future scene reconstruction is necessary for effective end-to-end planning. Through extensive experiments on the nuPlan benchmark (720 scenarios, over 11,000 samples), we show that task-driven semantic alignment—rather than pixel-level reconstruction—can achieve high planning accuracy while remaining efficient and interpretable. A lightweight planner operating on only 16 discrete intent tokens reaches 0.479 m ADE, demonstrating that strong performance can be achieved without dense reconstruction or auxiliary supervision, thus validating the effectiveness of our minimal, cognitively inspired design.
This work establishes a new paradigm in autonomous driving: planning through understanding. We demonstrate that:
1) High-quality inputs enable powerful yet simple planners. Reliable perception allows the planner to achieve strong performance without auxiliary reconstruction losses, suggesting that high accuracy does not require complex or over-parameterized models.
2) Future-aware decoding enables foresightful behavior. Conditioning trajectory generation on predicted intent tokens—rather than reconstructed scenes—yields a 12.6% improvement, illustrating that imagination-driven reasoning underlises intelligent control.
3) Explicit reconstruction loss offers no benefit. Explicit reconstruction loss offers no benefit; it constrains the model. Instead, a clean optimization signal focused solely on the final planning outcome naturally yields robust and generalizable intent representations.
The emergent property of ”temporal fuzziness”—the ability to focus on what matters semantically rather than rigidly aligning to when—further highlights the cognitive flexibility of our approach. This enables human-like adaptability under uncertainty.
Beyond autonomous driving, TIWM suggests a broader principle for embodied intelligence: true autonomy arises not from reconstructing the world in full fidelity, but from understanding its semantic essence and imagining possible futures. By unifying representation, intent, and prediction into a single, token-based framework, TIWM bridges the gap between data-driven perception and cognitively inspired reasoning. Future work will explore closed-loop evaluation, token scalability, and the generalization of belief–intent co-evolution to other domains. We view this as a step toward agents that plan through understanding, suggesting that sparse, task-aligned tokenization may represent a fundamental mechanism for efficient cognition in artificial systems.